OpenStudy (anonymous):

Blueprints of a house are drawn to scale of 1/4 in = 1 ft. A kitchen measures 3.5 in by 5 in. What is the actual size of the kitchen?

OpenStudy (anonymous):

I got 70 ft.

OpenStudy (anonymous):

Is that correct ?

hero (hero):

That's not what I got

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

What did you get,then?

OpenStudy (anonymous):

I multiplied 3.5 *5

hero (hero):

That's not what you were supposed to do.

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

Would you show me the procedure,then?

hero (hero):

The procedure is to solve proportions: \[\frac{.25}{3.5} = \frac{1}{x}\] and \[\frac{.25}{5} = \frac{1}{x}\]

OpenStudy (anonymous):

oh, you do them separately.

hero (hero):

After solving both those correctly for x, then you will have the dimensions in feet. Multiply them together.

hero (hero):

You don't have to do them separately. You can multiply together, but you can't forget to convert afterwards.

OpenStudy (anonymous):

Whatever you get, you multiply that together? both x's?

hero (hero):

Yes

OpenStudy (anonymous):

And that is the answer ? Ok. tnx!

OpenStudy (anonymous):

Let me do that.

OpenStudy (anonymous):

I get a long decimal for the first one.

OpenStudy (anonymous):

o. let me convert.